Pre-Final Course Grade – Chau, Thau

Lab/Discussion Average: 94.75 Exam Average: 84.9 (lowest score counts half as much as better two) Problem Set Average: 95.71 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 87.85 (based only on work to date)

Pre-Final Course Grade – Corrado, Alex

Lab/Discussion Average: 91.1 Exam Average: 61 (lowest score counts half as much as better two) Problem Set Average: 68.5 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 66.37 (based only on work to date)

Pre-Final Course Grade – Ferketic, Mary

Lab/Discussion Average: 94.75 Exam Average: 85.8 (lowest score counts half as much as better two) Problem Set Average: 76.11 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 85.7 (based only on work to date)

Pre-Final Course Grade – Fimognari, Allison

Lab/Discussion Average: 90.31 Exam Average: 94.6 (lowest score counts half as much as better two) Problem Set Average: 93.37 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 93.81 (based only on work to date)

Pre-Final Course Grade – Fournier, Rachel

Lab/Discussion Average: 96.75 Exam Average: 86.4 (lowest score counts half as much as better two) Problem Set Average: 90.47 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 88.46 (based only on work to date)

Pre-Final Course Grade – Gallant, Monique

Lab/Discussion Average: 94.93 Exam Average: 62.3 (lowest score counts half as much as better two) Problem Set Average: 80.01 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 69.49 (based only on work to date)

Pre-Final Course Grade – Hage, Brandon

Lab/Discussion Average: 90.64 Exam Average: 93.8 (lowest score counts half as much as better two) Problem Set Average: 84 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 91.95 (based only on work to date)

Pre-Final Course Grade – Hana, Brikena

Lab/Discussion Average: 90.64 Exam Average: 69.8 (lowest score counts half as much as better two) Problem Set Average: 85.65 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 75.04 (based only on work to date)

Pre-Final Course Grade – John, Stephanie

Lab/Discussion Average: 96.75 Exam Average: 62.8 (lowest score counts half as much as better two) Problem Set Average: 63.69 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 67.78 (based only on work to date)

Pre-Final Course Grade – Kico, Fioralda

Lab/Discussion Average: 96.82 Exam Average: 71.2 (lowest score counts half as much as better two) Problem Set Average: 66.21 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 74.15 (based only on work to date)

Pre-Final Course Grade – Kopp, Christopher

Lab/Discussion Average: 96.82 Exam Average: 95.6 (lowest score counts half as much as better two) Problem Set Average: 96.61 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 95.92 (based only on work to date)

Pre-Final Course Grade – Marsh, Michael

Lab/Discussion Average: 71.96 Exam Average: 84.6 (lowest score counts half as much as better two) Problem Set Average: 65.58 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 80.08 (based only on work to date)

Pre-Final Course Grade – Masury, Katherine

Lab/Discussion Average: 91.68 Exam Average: 74.1 (lowest score counts half as much as better two) Problem Set Average: 81.14 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 77.62 (based only on work to date)

Pre-Final Course Grade – McLaughlin, Tori

Lab/Discussion Average: 91.68 Exam Average: 74.6 (lowest score counts half as much as better two) Problem Set Average: 78.1 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 77.54 (based only on work to date)

Pre-Final Course Grade – Meng, Lingjie (Sarah)

Lab/Discussion Average: 83.37 Exam Average: 95.2 (lowest score counts half as much as better two) Problem Set Average: 92.65 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 93.15 (based only on work to date)

Pre-Final Course Grade – Ng, Janet

Lab/Discussion Average: 83.37 Exam Average: 81 (lowest score counts half as much as better two) Problem Set Average: 85.49 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 81.98 (based only on work to date)

Pre-Final Course Grade – Nguyen, Julie-Jennifer

Lab/Discussion Average: 91.89 Exam Average: 92.2 (lowest score counts half as much as better two) Problem Set Average: 89.73 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 91.8 (based only on work to date)

Pre-Final Course Grade – O’Brien, Kevin

Lab/Discussion Average: 91.89 Exam Average: 92.6 (lowest score counts half as much as better two) Problem Set Average: 70.3 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 89.3 (based only on work to date)

Pre-Final Course Grade – Oram, Thomas

Lab/Discussion Average: 78.85 Exam Average: 69.6 (lowest score counts half as much as better two) Problem Set Average: 84.12 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 72.99 (based only on work to date)

Pre-Final Course Grade – Seniuk, Keeley

Lab/Discussion Average: 81.84 Exam Average: 69.4 (lowest score counts half as much as better two) Problem Set Average: 81.73 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 72.94 (based only on work to date)

Pre-Final Course Grade – Smith, Amanda

Lab/Discussion Average: 81.84 Exam Average: 86.3 (lowest score counts half as much as better two) Problem Set Average: 77.27 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 84.37 (based only on work to date)

Pre-Final Course Grade – Torres, Sasha

Lab/Discussion Average: 96.25 Exam Average: 61.8 (lowest score counts half as much as better two) Problem Set Average: 57.86 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 66.16 (based only on work to date)

Pre-Final Course Grade – Tran, Thu

Lab/Discussion Average: 94.52 Exam Average: 80.4 (lowest score counts half as much as better two) Problem Set Average: 94.72 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 84.46 (based only on work to date)

Pre-Final Course Grade – Turpin, Julie

Lab/Discussion Average: 96.77 Exam Average: 83.2 (lowest score counts half as much as better two) Problem Set Average: 92.34 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 86.44 (based only on work to date)

Pre-Final Course Grade – Westin, Miriam

Lab/Discussion Average: 95.04 Exam Average: 93.6 (lowest score counts half as much as better two) Problem Set Average: 84.52 (These will go up in some cases in final

computation – I did not drop lowest 2 yet) Course Average: 92.51 (based only on work to date)